OpenStudy (anonymous):

how do you get the answer to (x-y=-5 (-x+4y=10

OpenStudy (anonymous):

all i know is you use substitution

OpenStudy (anonymous):

add the two equations and post what you got.

OpenStudy (anonymous):

what do you mean by add them

OpenStudy (anonymous):

steps: x-y=-5 -x+4y=10 despejas x x=-5+y 5-y+4y=10 3y=5 y=5/3 ahora en x-y=-5 x-(5/3)=-5 x=-10/3

OpenStudy (anonymous):

x-y = 5 -------> Equation 1 -x+4y = 10 ------------> Equation 2 Add equations 1 and 2

OpenStudy (anonymous):

oh also if it helps im only in 8th grade

OpenStudy (shadowfiend):

By `add' he means basically add each part (the parts with x, the parts with y, and the parts with neither) in each equation, and see what the resulting equation is.

OpenStudy (anonymous):

No shadow fiend. The problem is different. These answers were getting jumbled up, with another post

OpenStudy (anonymous):

so -x=4y=5

OpenStudy (anonymous):

if (3 x 4 ) = 12 and (5 x 4) = 20, then when we add the equations we get (3 x 4) + (5 x 4) = 12 + 20 12 + 20 = 32 32 = 32. so it checks out. similarly add your equations,

OpenStudy (anonymous):

meaning (x-y) +(-x+4y) = -5 + 10

OpenStudy (anonymous):

wow this is even more confusing now

OpenStudy (anonymous):

what is confusing about this?

OpenStudy (shadowfiend):

So, norbeybarajas gave another good answer, both will lead you to the same path. If you add the two equations up, you'll see that the x disappears and you can solve for y.

OpenStudy (anonymous):

ok thx for the answers not confusing any more

OpenStudy (anonymous):

do you understand that if a = b and c = d, then a+c will be equal to b+d?

OpenStudy (anonymous):

yeah

OpenStudy (shadowfiend):

Rock it. Make sure you click `good answer' next to the answers that were helpful :)

OpenStudy (anonymous):

alright can do

OpenStudy (anonymous):

okay, so that is exactly what I did in your equation. I saw that one equation had x and the other equation had -x. so I knew that if i added them up, the x will cancel each other out, leaving me only with y. which means i can solve for y.

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

once I solve for y, all i need to do is substitute it in one of the equations and get x.